OpenStudy (nuccioreggie):

There are (64)2 ⋅ 60 pigs on a farm. What is the total number of pigs on the farm? 0 66 68 648

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

hint: \[\begin{gathered} {6^0} = 1 \hfill \\ {\left( {{6^4}} \right)^2} = {6^{4 \times 2}} = {6^8} \hfill \\ \end{gathered} \]

OpenStudy (nuccioreggie):

C

OpenStudy (michele_laino):

that's right!